The adlslides package^{∗}

Andrew D. Lewis

1998/07/28

1 The adlslides package

The package adlslides provides some macros for making what I think are nice slides. It is built on top of Timothy Van Zandt’s excellent seminar package. This is, you need to have this to use adlslides. You also need his pstricks package which is loaded by adlcolors.

2 Goodies provided by adlslides

First lets describe the options.

1. The option pson says that you have a file mypsfonts.sty available which specifies the PostScript font you will be using for your talk. If you specifiy this, I will assume you know what you are doing.

2. There is also psoff…

3. If you want color slides, use the color option.

4. If you want a colored background, use the backgroundcolor option.

5. If you want particularly snazzy borders for your slides, use the fancyframes option. This must appear after the color option if the latter is specified.

6. If you want the snazzy borders, but no colour in other places in your slides, use the option simplecolor.

7. There are options for the various places you have worked. For me these are caltech, warwick, and queens. You need to make these so they make sense for you. Depending on which you choose, the logo on the title page will vary, and the information on the headings of the slides will differ. You may also choose plain which gives no headers.

∗This file has version number v1.0b, last revised 1998/07/28.

Also on the topic of colours, the package adlcolors is loaded (it is included in the distribution). At the end of this file you will find various default settings for colours used in slides (section headers, list items, etc.). Feel free to change these.

Sometimes I need to have certain things in the document depend on whether

\colortest I am using colour or not. I do this with the \colortest command which is used

thusly: \colortest{command if I am using colour}{command if I am not using

\psfonttest colour}. There is a similar command \psfonttest for PostScript fonts. \maketitleslide For making the title slide, use the \maketitleslide command. This in-

\maketitle vokes the usual \maketitle command, so before you use \maketitleslide you should have all the pertinent information entered (title, author, date). The \maketitleslide command is compatible with the \maketitle command from my adlart package. Thus you can use both adlslides and adlart. In the seminar manner, \maketitleslide will make a landscape title slide, and

\maketitleslide*  \maketitleslide* will make a portrait title slide.

\outline You can also make an automated outline slide using the \outline command. Unlike the \maketitleslide command, \outline must be manually placed in a slide environment. That is to say, it is not a “\outlineslide” command, if you catch my meaning.

It is nice sometimes to make copies of your talk to give to people. When you do this, you may want to include explanatory notes so things on the slides will make sense to someone who does not remember the talk. To facilitate this, I use the

notes notes environment. To use this, first end the current slide with \end{slide} (this

is a seminar thing to do), and then write your note within the notes environment. Equations are numbered globally, not by section. I use the seminar package to create pdf’s to use with data projectors. In this

case, one often wishes to incrementally display the page. This was designed to be done in seminar with printed transparencies. To do this for projection, I have

\overlaystrue changed the overlay method for seminar. To invoke this, use the \overlaystrue \overlaysfalse command. Turn this off with the \overlaysfalse command. I also have made a

footnote command which will display the footnotes incrementally. The command

\custfootnote is \custfootnote[n]{num}{text}. Here n is the overlay where the footnote

should appear, num is the footnote number, and text is the text of the footnote.

There are various style hacks I have made, but I hope these will be transparent to you, so “good” are they. They involve things like changing the indentation in list environments.

I have included a sample file which should get you started on writing your talk.

3 Implementation

You will have to make some changes here for this all to work! The changes that need to be made are all to do with where you work, and will specify the logo used on the title slide, and your affiliation which is on the top of every slide. It should be easy to see how to make the necessary modifications.

3.1 adlslides.sty

1 ⟨∗adlslides⟩

Load the necessary packages, declare options, etc. There’s an annoying kludge here when loading fancybox.sty which redefines \c@tocdepth. It needs to be reset to 3…

2

 3 \RequirePackage{pstcol,removefromreset,graphicx}
 4 \newif\if@ADL@color\@ADL@colorfalse
 5 \newif\if@ADL@simplecolor\@ADL@simplecolorfalse
 6 \DeclareOption{pson}{\def\@ADL@@ps@on{\@ne}}
 7 \DeclareOption{psoff}{\def\@ADL@@ps@on{\z@}}
 8 \DeclareOption{color}{\@ADL@colortrue}
 9 \DeclareOption{simplecolor}{\@ADL@simplecolortrue\@ADL@colortrue}
10 \DeclareOption{fancyframes}{\if@ADL@color
11 \AtBeginDocument{\makeatletter\input{semcolor.sty}
12 \makeatother\slideframe[\psset{linecolor=red,shadowsize=6pt}]{scshadow}}
13 \else\AtBeginDocument{\makeatletter\input{fancybox.sty}\c@tocdepth\thr@@
14 \makeatother\slideframe{Oval}}\fi}
15 \DeclareOption{backgroundcolor}{\@ADL@colortrue\AtBeginDocument{%
16 \makeatletter\input{semcolor.sty}%
17 \let\@ADL@old@title\@title%
18 \gdef\@title{\seccolor\@ADL@old@title}
19 \slideframe{none}\pagecolor{background}\foreground}}
20 \DeclareOption{caltech}{\def\@ADL@school{caltech}}
21 \DeclareOption{queens}{\def\@ADL@school{queens}}
22 \DeclareOption{department}{\def\@ADL@school{department}}
23 \DeclareOption{mathengg}{\def\@ADL@school{mathengg}}
24 \DeclareOption{plain}{\def\@ADL@school{plain}}
25 \DeclareOption{warwick}{\def\@ADL@school{warwick}}
26 \DeclareOption{apsc}{\def\@ADL@school{apsc}}
27 \ExecuteOptions{psoff,plain}
28 \ProcessOptions
29 \if@ADL@simplecolor\@ADL@colorfalse\fi

Undo what I think is some silliness done by seminar.

30

31 \renewcommand\normalsize{%
32 \@setfontsize\normalsize\@xpt\@xiipt
33 \abovedisplayskip 10\p@ \@plus2\p@ \@minus5\p@
34 \abovedisplayshortskip \z@ \@plus3\p@
35 \belowdisplayshortskip 6\p@ \@plus3\p@ \@minus3\p@
36 \belowdisplayskip \abovedisplayskip
37 \let\@listi\@listI}

Do some fiddling with the PostScript fonts stuff. While we’re defining tests, do the colortest as well.

38

39 \ifnum\@ADL@@ps@on=\@ne\RequirePackage{mypsfont}\fi
40 \@ifundefined{psfonttest}{%

41 \newcommand\psfonttest[2]{\ifnum\@ADL@@ps@on=\@ne#1\else#2\fi}}{}
42 \newcommand\colortest[2]{\if@ADL@color#1\else#2\fi}

Here is where you define the page styles for all those places which have had the honour of having you as an employee.

43

44 \newpagestyle{mypage}{}{}
45 \def\@ADL@caltech{caltech}
46 \def\@ADL@plain{plain}
47 \def\@ADL@department{department}
48 \def\@ADL@mathengg{mathengg}
49 \def\@ADL@queens{queens}
50 \def\@ADL@warwick{warwick}

51

52 \ifx\@ADL@school\@ADL@caltech
53 \renewpagestyle{mypage}{}
54 {\tiny\foreground Andrew D.~Lewis\hfill California Institute of Technology}
55 \else
56 \ifx\@ADL@school\@ADL@warwick
57 \renewpagestyle{mypage}{}
58 {\tiny\foreground Andrew D.~Lewis\hfill University of Warwick}
59 \else
60 \ifx\@ADL@school\@ADL@queens
61 \renewpagestyle{mypage}{}
62 {\tiny\foreground Andrew D.~Lewis\hfill Queen's University, Mathematics \& Statistics}
63 \else
64 \ifx\@ADL@school\@ADL@department
65 \renewpagestyle{mypage}{}
66 {\tiny\foreground Mathematics \& Statistics\hfill Queen's University}
67 \ifx\@ADL@school\@ADL@mathengg
68 \renewpagestyle{mypage}{}
69 {\tiny\foreground Mathematics \& Engineering\hfill Queen's University}
70 \fi\fi\fi\fi\fi
71 \pagestyle{mypage}

If you are using colour, then modify the list environments appropriately. Even if we are not using colour, the standard lists need to be modified for indentation. So I define \slidelistenv environments.

72

73 \if@ADL@color
74 \def\slideenumerate{%
75 \ifnum \@enumdepth >\thr@@\@toodeep\else
76  \advance\@enumdepth\@ne
77  \edef\@enumctr{enum\romannumeral\the\@enumdepth}%
78  \expandafter
79  \list
80  \csname label\@enumctr\endcsname
81  {\usecounter\@enumctr%
82  \def\makelabel##1{\hss\llap{\enumcolor##1}}}%
83 \fi}
84 \else

 85 \let\slideenumerate\enumerate
 86 \fi
 87 \if@ADL@color
 88 \@ifundefined{proclist}%
 89 {\PackageWarning{adlslides}{If using ``proclist'' package, load first.}}{}
 90 \def\slideproclist{\ifnum \@ADL@plistdepth >3 \@toodeep\else
 91 \advance\@ADL@plistdepth \@ne
 92 \edef\@ADL@plistctr{plist\romannumeral\the\@ADL@plistdepth}\list
 93 {\csname label\@ADL@plistctr\endcsname}{\usecounter{\@ADL@plistctr}
 94 \topsep 3\p@ plus \p@ minus \p@\parsep 1.5\p@ plus \p@ minus .5\p@
 95 \itemsep\parsep\partopsep \p@ plus \p@ minus \p@
 96 \def\makelabel##1{\hss\llap{\enumcolor##1}}}\fi}
 97 {\endlist}
 98 \else
 99 \let\slideproclist\proclist
100 \fi
101 \if@ADL@color
102 \def\slideitemize{\itemindent\z@\settowidth{\labelwidth}{\labelitemi}%
103 \leftmargini\labelwidth\advance\leftmargini\labelsep%
104 \ifnum \@itemdepth >\thr@@\@toodeep\else
105  \advance\@itemdepth\@ne
106  \edef\@itemitem{labelitem\romannumeral\the\@itemdepth}%
107  \expandafter
108  \list
109  \csname\@itemitem\endcsname
110  {\def\makelabel##1{\hss\llap{\itemcolor##1}}}%
111 \fi}
112 \else
113 \def\slideitemize{\itemindent\z@\settowidth{\labelwidth}{\labelitemi}%
114 \leftmargini\labelwidth\advance\leftmargini\labelsep%
115 \ifnum \@itemdepth >\thr@@\@toodeep\else
116  \advance\@itemdepth\@ne
117  \edef\@itemitem{labelitem\romannumeral\the\@itemdepth}%
118  \expandafter
119  \list
120  \csname\@itemitem\endcsname
121  {\def\makelabel##1{\hss\llap{##1}}}%
122 \fi}
123 \fi
124 \if@ADL@color
125 \newcommand*\slidedescriptionlabel[1]{\hspace\labelsep
126 \descriptioncolor\normalfont\bfseries #1}
127 \else
128 \let\slidedescriptionlabel\descriptionlabel
129 \fi

If colour section commands are not already provided (for example by adlart), define them.

130

131 \if@ADL@color

132 \@ifundefined{sectioncolor}{%
133 \let\intsection\section
134 \def\sectioncolor{\@ifstar{\@sectioncolorstar}{\@sectioncolornostar}}
135 \def\@sectioncolorstar#1{\intsection*{\seccolor#1}}
136 \def\@sectioncolornostar#1{\intsection{\seccolor#1}}
137 \let\intsubsection\subsection
138 \def\subsectioncolor{\@ifstar{\@subsectioncolorstar}
139 {\@subsectioncolornostar}}
140 \def\@subsectioncolorstar#1{\intsubsection*{\seccolor#1}}
141 \def\@subsectioncolornostar#1{\intsubsection{\seccolor#1}}
142 }{}
143 \fi

If you are using colour, then modify the slide commands so that the colour list environments and colour sections are used. If colours are not used, then we still need to modify the list environments within the slides.

144

145 \if@ADL@color
146 \def\slide{\let\itemize\slideitemize\let\enumerate\slideenumerate%
147 \let\descriptionlabel\slidedescriptionlabel\let\proclist\slideproclist%
148 \let\@ADL@emph\emph\renewcommand\emph[1]{\@ADL@emph{\emphcolor##1}}%
149 \@ifundefined{cfigure}{}{%
150 \let\@ADLtemp@cfigure\cfigure
151 \renewcommand\cfigure[1]{\@ADLtemp@cfigure{##1color}}%
152 \let\@ADLtemp@ctwofigure\ctwofigure\renewcommand\ctwofigure[2]{%
153 \@ADLtemp@ctwofigure{##1color}{##2color}}}
154 \let\section\sectioncolor\let\subsection\subsectioncolor%
155 \NestedSlide@Error{slide}%
156 \landscapetrue\foreground
157 \@ifnextchar[{\begin@slide}{\begin@slide[\slidewidth,\slideheight]}}
158 \@namedef{slide*}{\let\itemize\slideitemize\let\enumerate\slideenumerate%
159 \let\descriptionlabel\slidedescriptionlabel\let\proclist\slideproclist%
160 \let\@ADL@emph\emph\renewcommand\emph[1]{\@ADL@emph{\emphcolor##1}}%
161 \@ifundefined{cfigure}{}{%
162 \let\@ADLtemp@cfigure\cfigure
163 \renewcommand\cfigure[1]{\@ADLtemp@cfigure{##1color}}%
164 \let\@ADLtemp@ctwofigure\ctwofigure\renewcommand\ctwofigure[2]{%
165 \@ADLtemp@ctwofigure{##1color}{##2color}}}
166 \let\section\sectioncolor\let\subsection\subsectioncolor%
167 \NestedSlide@Error{slide*}%
168 \landscapefalse\foreground
169 \@ifnextchar[{\begin@slide}{\begin@slide[\slidewidth,\slideheight]}}
170 \else
171 \def\slide{\let\itemize\slideitemize\let\enumerate\slideenumerate%
172 \let\descriptionlabel\slidedescriptionlabel\let\proclist\slideproclist%
173 \NestedSlide@Error{slide}%
174 \landscapetrue\foreground
175 \@ifnextchar[{\begin@slide}{\begin@slide[\slidewidth,\slideheight]}}
176 \@namedef{slide*}{\let\itemize\slideitemize\let\enumerate\slideenumerate%
177 \let\descriptionlabel\slidedescriptionlabel\let\proclist\slideproclist%

178 \NestedSlide@Error{slide*}%
179 \landscapefalse\foreground
180 \@ifnextchar[{\begin@slide}{\begin@slide[\slidewidth,\slideheight]}}
181 \fi

Macros for building the title slide. The logo modifications here will have to be user specified.

182

183 \newcommand\maketitleslide{\@ifstar{\@ADL@maketitleslide}
184 {\@ADL@@maketitleslide}}
185 \newcommand\@ADL@maketitleslide{{\pagestyle{empty}
186 \begin{slide*}\global\c@slide\z@
187 \maketitle\vspace{0.15in}
188 \ifx\@ADL@school\@ADL@caltech
189 \begin{center}%
190 \includegraphics[width=1in]{/home/andrew/lib/texmf/tex/logos/citlogo}%
191 \end{center}\fi%
192 \ifx\@ADL@school\@ADL@warwick
193 \begin{center}%
194 \includegraphics[height=1in]{/home/andrew/lib/texmf/tex/logos/warwicklogo}%
195 \end{center}\fi%
196 \ifx\@ADL@school\@ADL@queens
197 \begin{center}%
198 \includegraphics[width=0.8in]{/home/andrew/lib/texmf/tex/logos/queenslogo}%
199 \end{center}%
200 \fi\vspace{0.2in}\end{slide*}}}
201 \newcommand\@ADL@@maketitleslide{{\pagestyle{empty}
202 \begin{slide}\global\c@slide\z@
203 \maketitle\ifx\@ADL@school\@ADL@caltech
204 \begin{center}%
205 \includegraphics[width=0.8in]{/home/andrew/lib/texmf/tex/logos/citlogo}%
206 \end{center}\fi%
207 \ifx\@ADL@school\@ADL@warwick
208 \begin{center}%
209 \includegraphics[height=0.8in]{/home/andrew/lib/texmf/tex/logos/warwicklogo}%
210 \end{center}\fi%
211 \ifx\@ADL@school\@ADL@queens
212 \begin{center}%
213 \includegraphics[width=0.8in]{/home/andrew/lib/texmf/tex/logos/queenslogo}%
214 \end{center}%
215 \fi\vspace{0.1in}\end{slide}}}

A usual \maketitle command in the event you do not want to use the \maketitleslide command with all its constraints. This is also used internally by \maketitleslide. Also get rid of that pesky \slidemaketitle defined by

 seminar. . .

216

217 \renewcommand\maketitle{\par
218 \begingroup
219  \renewcommand\thefootnote{\fnsymbol{footnote}}%

220  \def\@makefnmark{\hbox to\z@{$\m@th^{\@thefnmark}$\hss}}%
221  \long\def\@makefntext##1{\parindent 1em\noindent
222  \hbox to0.3em{\hss$\m@th^{\@thefnmark}$}\hangindent0.3em##1}%
223  \if@twocolumn
224  \ifnum \col@number=\@ne
225  \@maketitle
226  \else
227  \twocolumn[\@maketitle]%
228  \fi
229  \else
230  \newpage
231  \global\@topnum\z@   % Prevents figures from going at top of page.
232  \@maketitle
233  \fi\@thanks
234 \endgroup
235 \setcounter{footnote}{0}%
236 \let\@ADL@intthanks\relax
237 \let\maketitle\relax\let\@maketitle\relax
238 \gdef\@thanks{}\gdef\@author{}\gdef\@title{}}
239 \let\slidemaketitle\maketitle

For making the outline slide.

240

241 \@ifundefined{outline}{
242 \newcommand\outline{\section*{Outline}\@starttoc{toc}}}
243 {\renewcommand\outline{\section*{Outline}\@starttoc{toc}}}

244

245 \renewcommand\l@section[2]{%
246 \ifnum \c@tocdepth >\z@
247  \addpenalty{\@secpenalty}%
248  \addvspace{0.5em \@plus\p@}%
249  \setlength\@tempdima{1.5em}%
250  \begingroup
251  \parindent \z@ \rightskip \@pnumwidth
252  \parfillskip -\@pnumwidth
253  \leavevmode
254  \advance\leftskip\@tempdima
255  \hskip -\leftskip
256  #1\nobreak\hfil \nobreak\hbox to\@pnumwidth{\hss \phantom{#2}}\par
257  \endgroup
258 \fi}
259 \def\@dottedtocline#1#2#3#4#5{\relax}

Some style preferences: equations numbered globally, and use a hollow circle the bullet for second level itemised lists (I had problems confusing the dashes for minus signs).

260

261 \renewcommand\theequation{\arabic{equation}}
262 \@removefromreset{equation}{section}
263 \renewcommand\labelitemii{$\circ$}
264 \slidewidth 8.5in\slideheight 6.3in

If we are only writing slides, then ignore the notes by making them comments with the comment package. Otherwise, define the notes environment.

265 \ifslidesonly
266 \RequirePackage{comment}\excludecomment{notes}
267 \else
268 \newenvironment{notes}
269 {\null\vfil\par\noindent{\large\bfseries Notes for Slide \theslide}%
270 \hfill\break\vspace{0.0in}\par\noindent\ignorespaces}
271 {\vfil\newpage}
272 \fi

Make footnotes numbered.

273

274 \renewcommand\theslidefootnote{\arabic{footnote}}

275

276 \definecolor{section}{cmyk}{0,0.89,0.94,0.28}
277 \definecolor{border}{cmyk}{0,0.89,0.94,0.28}
278 \definecolor{emphasise}{cmyk}{0.92,0,0.59,0.25}
279 %\definecolor{background}{rgb}{0,0,1}
280 \definecolor{background}{rgb}{1,1,1}
281 \if@ADL@color
282 \newcommand\foreground{\color{black}\sffamily}
283 \else
284 \newcommand\foreground{\relax}
285 \fi
286 \newcommand\seccolor{\color{section}\sffamily}
287 \newcommand\itemcolor{\color{red}}
288 \newcommand\enumcolor{\color{red}}
289 \newcommand\descriptioncolor{\color{red}}
290 \newcommand\emphcolor{\color{emphasise}}

291

Make it so that overlays add to the slide.

292

293 \def\pst@initoverlay#1{%
294 \pst@Verb{%
295 /BeginOL {dup (all) eq exch TheOL le or {IfVisible not {Visible
296 /IfVisible true def} if} {IfVisible {Invisible /IfVisible false def} if}
297 ifelse} def
298 \tx@InitOL /TheOL (#1) def}}

299

Add command so that footnotes work okay with overlays.

300

301 \def\custfootnote{\@ifnextchar[\@ADL@custfootnoteol\@ADL@custfootnote}
302 \def\@ADL@custfootnote#1#2{%
303 \begin{overlay}{0}\@ADL@custfootnotemark{#1}\end{overlay}%
304 \@ADL@custfootnotetext{0}{#2}}
305 \def\@ADL@custfootnoteol[#1]#2#3{\begin{overlay}{#1}\@ADL@custfootnotemark{#2}%
306 \end{overlay}\@ADL@custfootnotetext{#1}{#2}{#3}}
307 \long\def\@ADL@custfootnotetext#1#2#3{\begin{overlay}{#1}\insert\footins{%

308  \reset@font\footnotesize%
309  \interlinepenalty\interfootnotelinepenalty%
310  \splittopskip\footnotesep%
311  \splitmaxdepth\dp\strutbox\floatingpenalty\@MM%
312  \hsize\columnwidth\@parboxrestore%
313  \color@begingroup%
314  \@ADL@custmakefntext{#1}{#2}{%
315  \rule\z@\footnotesep\ignorespaces#3\@finalstrut\strutbox}%
316  \color@endgroup}\end{overlay}}
317 \long\def\@ADL@custmakefntext#1#2#3{\begin{overlay}{#1}\parindent 1em\noindent%
318  \hb@xt@1.8em{\hss$\@textsuperscript{\normalfont$#2$}$}#3%
319  \end{overlay}}
320 \def\@ADL@custfootnotemark#1{%
321 \leavevmode%
322 \ifhmode\edef\@x@sf{\the\spacefactor}\nobreak\fi%
323 \@ADL@custmakefnmark{#1}%
324 \ifhmode\spacefactor\@x@sf\fi%
325 \relax}
326 \def\@ADL@custmakefnmark#1{\hbox{\@textsuperscript{\normalfont$#1$}}}

327

328 ⟨/adlslides⟩

3.2 adlcolors.sty

This is easy…

329 ⟨∗adlcolors⟩

330

331 \RequirePackage{pstcol}

332

333 \definecolor{greenyellow}{cmyk}{0.15,0,0.69,0}
334 \definecolor{goldenrod}{cmyk}{0,0.10,0.84,0}
335 \definecolor{dandelion}{cmyk}{0,0.29,0.84,0}
336 \definecolor{apricot}{cmyk}{0,0.32,0.52,0}
337 \definecolor{peach}{cmyk}{0,0.50,0.70,0}
338 \definecolor{melon}{cmyk}{0,0.46,0.50,0}
339 \definecolor{yelloworange}{cmyk}{0,0.42,1,0}
340 \definecolor{orange}{cmyk}{0,0.61,0.87,0}
341 \definecolor{burntorange}{cmyk}{0,0.51,1,0}
342 \definecolor{bittersweet}{cmyk}{0,0.75,1,0.24}
343 \definecolor{redorange}{cmyk}{0,0.77,0.87,0}
344 \definecolor{mahogany}{cmyk}{0,0.85,0.87,0.35}
345 \definecolor{maroon}{cmyk}{0,0.87,0.68,0.32}
346 \definecolor{brickred}{cmyk}{0,0.89,0.94,0.28}
347 \definecolor{orangered}{cmyk}{0,1,0.50,0}
348 \definecolor{rubinered}{cmyk}{0,1,0.13,0}
349 \definecolor{wildstrawberry}{cmyk}{0,0.96,0.39,0}
350 \definecolor{salmon}{cmyk}{0,0.53,0.38,0}
351 \definecolor{carnationpink}{cmyk}{0,0.63,0,0}
352 \definecolor{violetred}{cmyk}{0,0.81,0,0}

353 \definecolor{rhodamine}{cmyk}{0,0.82,0,0}
354 \definecolor{mulberry}{cmyk}{0.34,0.90,0,0.02}
355 \definecolor{redviolet}{cmyk}{0.07,0.90,0,0.34}
356 \definecolor{fuchsia}{cmyk}{0.47,0.91,0,0.08}
357 \definecolor{lavender}{cmyk}{0,0.48,0,0}
358 \definecolor{thistle}{cmyk}{0.12,0.59,0,0}
359 \definecolor{orchid}{cmyk}{0.32,0.64,0,0}
360 \definecolor{darkorchid}{cmyk}{0.40,0.80,0.20,0}
361 \definecolor{purple}{cmyk}{0.45,0.86,0,0}
362 \definecolor{plum}{cmyk}{0.50,1,0,0}
363 \definecolor{violet}{cmyk}{0.79,0.88,0,0}
364 \definecolor{royalpurple}{cmyk}{0.75,0.90,0,0}
365 \definecolor{blueviolet}{cmyk}{0.86,0.91,0,0.04}
366 \definecolor{periwinkle}{cmyk}{0.57,0.55,0,0}
367 \definecolor{cadetblue}{cmyk}{0.62,0.57,0.23,0}
368 \definecolor{cornflowerblue}{cmyk}{0.65,0.13,0,0}
369 \definecolor{midnightblue}{cmyk}{0.98,0.13,0,0.43}
370 \definecolor{navyblue}{cmyk}{0.94,0.54,0,0}
371 \definecolor{royalblue}{cmyk}{1,0.50,0,0}
372 \definecolor{cerulean}{cmyk}{0.94,0.11,0,0}
373 \definecolor{processblue}{cmyk}{0.96,0,0,0}
374 \definecolor{skyblue}{cmyk}{0.62,0,0.12,0}
375 \definecolor{turquoise}{cmyk}{0.85,0,0.20,0}
376 \definecolor{tealblue}{cmyk}{0.86,0,0.34,0.02}
377 \definecolor{aquamarine}{cmyk}{0.82,0,0.30,0}
378 \definecolor{bluegreen}{cmyk}{0.85,0,0.33,0}
379 \definecolor{emerald}{cmyk}{1,0,0.50,0}
380 \definecolor{junglegreen}{cmyk}{0.99,0,0.52,0}
381 \definecolor{seagreen}{cmyk}{0.69,0,0.50,0}
382 \definecolor{forestgreen}{cmyk}{0.91,0,0.88,0.12}
383 \definecolor{pinegreen}{cmyk}{0.92,0,0.59,0.25}
384 \definecolor{limegreen}{cmyk}{0.50,0,1,0}
385 \definecolor{yellowgreen}{cmyk}{0.44,0,0.74,0}
386 \definecolor{springgreen}{cmyk}{0.26,0,0.76,0}
387 \definecolor{olivegreen}{cmyk}{0.64,0,0.95,0.40}
388 \definecolor{rawsienna}{cmyk}{0,0.72,1,0.45}
389 \definecolor{sepia}{cmyk}{0,0.83,1,0.70}
390 \definecolor{brown}{cmyk}{0,0.81,1,0.60}
391 \definecolor{tan}{cmyk}{0.14,0.42,0.56,0}

392

393 ⟨/adlcolors⟩